# Online Resource Allocation via Static Bundle Pricing

### v1

@@ 7 Lower Bounds
We now construct the lower-bound instance in the prophet-inequality setting. Let $m\geq 16$ denote the number of items and suppose that we have exactly $B$ copies of each item. Set
$t=t_{m,B}\coloneqq\left\lfloor\left(\frac{m}{2B\ln m}\right)^{1/(B+2)}\right\rfloor,$
and let $N=Bt^{t}\leq g(m,t,B+1)$. By Lemma 7.3, there exists a $(B+1)$-way qualitatively independent family
$\mathcal{F}=\left\{\mathcal{P}^{(1)},\dots,\mathcal{P}^{(N)}\right\}$
of $N$ partitions of the $m$ items into $t$ classes each:
$\mathcal{P}^{(i)}=\bigl\{P^{(i)}_{1},\dots,P^{(i)}_{t}\bigr\},\qquad i\in[N].$
For each partition $i\in[N]$ and each class $a\in[t]$, introduce a single-minded buyer $B_{i,a}$ who wants exactly the bundle $P^{(i)}_{a}$ and whose value is
independently over all pairs $(i,a)$.
@@
For a fixed $i$, the number of buyers in the $i$-th group $\{B_{i,1},\dots,B_{i,t}\}$ whose value is $1$ is distributed as $\textnormal{Bin}(t,1/t)$. For each group $i\in[N]$, let $E_{i}$ be the event that all $t$ buyers in group $i$ have value $1$. These events are independent and satisfy
$\mathbb{P}\!\left[E_{i}\right]=p=(1/t)^{t}.$
Let $X$ be the number of groups in which all $t$ buyers have value $1$; then
$X\sim\textnormal{Bin}\bigl(N,(1/t)^{t}\bigr),\qquad\mathbb{E}\left[X\right]=Np=B.$
By Lemma G.3, we get $\mathbb{P}\!\left[X\geq B\right]\geq 0.5$. We next relate the combinatorial structure of the partitions to the capacity constraints via to simple observations.
Observation 7.4.
In any feasible allocation (respecting the $B$-copy constraint for each item), it is impossible to serve buyers from more than $B$ distinct groups.
Observation 7.5.
Suppose $E_{i_{1}},\dots,E_{i_{B}}$ all occur for $B$ distinct groups $i_{1},\dots,i_{B}$. Then there exists a feasible allocation that serves all $tB$ buyers in these $B$ groups.
We now bound the offline and online performance.
Offline Prophet.
Whenever $X\geq B$, by Observation 7.5 the offline optimum can choose any $B$ groups $i_{1},\dots,i_{B}$ among those with $E_{i_{j}}$ and serve all $t$ buyers in each of them, obtaining value $Bt$. Thus
$\mathrm{OPT}\;\geq\;Bt\cdot\mathbbm{1}{\left\{\{X\geq B\}\right\}},$
and hence
$\mathbb{E}\left[\mathrm{OPT}\right]\;\geq\;Bt\cdot\mathbb{P}\!\left[X\geq B\right]\;\geq\;0.5\,Bt$
for all sufficiently large $m$.
Online Algorithm.
Consider an arbitrary online algorithm. For each group $i\in[N]$, let $Y_{i}$ be the number of buyers in group $i$ with value $1$. Then $Y_{i}\sim\textnormal{Bin}(t,1/t)$ and thus $\mathbb{E}\left[Y_{i}\right]=1$. By Observation 7.4, the algorithm can serve buyers from at most $B$ distinct groups. Let $t_{i}$ denote the $i$-th different group (chronologically) that the algorithm selects to serve, thus:
$\mathbb{E}\left[\mathrm{ALG}\right]\;\leq\;\sum_{i=1}^{B}\mathbb{E}\left[\text{value from group $t_{i}$}\right]\;\leq\;2B.$
since after we select a buyer from group $t_{i}$ the expected total value of the remaining buyers from this group is less than $1$.
Competitive ratio.
Combining the two bounds, for all sufficiently large $m$,
$\frac{\mathbb{E}[\mathrm{OPT}]}{\mathbb{E}[\mathrm{ALG}]}\;\geq\;\frac{0.5\,Bt}{2B}=0.25\,t.$
Recall that
$t=\left\lfloor\left(\frac{m}{2B\ln m}\right)^{1/(B+2)}\right\rfloor$
For $m\geq 16$, the term in the floor is at least $1$, so
$t\;\geq\;\frac{1}{2}\left(\frac{m}{2B\ln m}\right)^{1/(B+2)}\;=\;\frac{1}{2}\,\left(\frac{1}{2B}\right)^{1/(B+2)}\left(\frac{m}{\ln m}\right)^{1/(B+2)}\geq\frac{1}{2}\,\cdot\frac{1}{2}\cdot\left(\frac{m}{\ln m}\right)^{1/(B+2)}.$
Thus:
$\frac{\mathbb{E}[\mathrm{OPT}]}{\mathbb{E}[\mathrm{ALG}]}\;\geq\;\frac{1}{16}\left(\frac{m}{\ln m}\right)^{1/(B+2)}.$
which proves Theorem˜1.4.
7.2 Proof of Theorem 1.5
We give a variant of the lower-bound template from Theorem˜1.4 that additionally enforces a maximum desired bundle size $d$. First we introduce the notion of balanced $r$-way QI partitions.
Balanced r-way QI Partitions.
We now refine the probabilistic argument of [poljak] to the balanced setting. The only difference is that we sample partitions from uniformly random balanced labelings.
Lemma 7.6 (Balanced refinement of Theorem 4 in [poljak]).
Let $n,t,r\geq 1$ be integers with $t\mid n$, and set $d:=n/t$. Then
$g_{\mathrm{bal}}(n,t,r,d)\;\geq\;\frac{e\,r}{t}\cdot\exp\!\left(\frac{n}{rt^{r}}\right),$
Proof.
For each $u\in U$ and each $\ell\in[r]$,
$\mathbb{P}\!\left[u\in P^{(j_{\ell})}_{a_{\ell}}\right]=\frac{d}{n}=\frac{1}{t},$
and by independence across the $r$ sampled partitions,
Hence $\mathbb{E}\left[Z\right]=\frac{n}{t^{r}}$. Moreover, for a fixed partition $j$ and class $a$, $P^{(j)}_{a}$ is a uniformly random subset of $U$ of size $d$. It is a well known fact that for a uniformly random size-$d$ subset, the membership indicators are negatively associated (sampling without replacement) therefore
Similarly to [poljak], choosing
$k=\left\lceil\frac{er}{t}\cdot\exp\!\left(\frac{n}{rt^{r}}\right)\right\rceil$
We now set $r:=B+1$, since $(B+1)$-way QI is exactly what rules out serving buyers from more than $B$ groups when there are $B$ copies of each item (as in Observation 7.4). We will take $m=dt$ items and require that every class in our partitions has size at most $d$; with $t\mid m$, balanced sampling gives classes of size exactly $m/t=d$.
Define, for $d$ sufficiently large and $1\leq B<\ln d$,
In Appendix G we prove the following.
Lemma 7.7.
With $t$ and $m$ as in (1),
$g_{\mathrm{bal}}(m,t,B+1,d)\;\geq\;B\,t^{\,t},$
for all sufficiently large $d$.
Lower-Bound Instance with Bundle Cap d.
Let $m=dt$ items and suppose there are exactly $B$ copies of each item. By Lemma˜7.7, there exists a $(B+1)$-way QI family
of $d$-balanced partitions of the $m$ items into $t$ classes each:
$\mathcal{P}^{(i)}=\{P^{(i)}_{1},\dots,P^{(i)}_{t}\},\qquad|P^{(i)}_{a}|\leq d.$
Analysis.
Define events $E_{i}$ and the random variable $X$ exactly as in Theorem˜1.4. Then $X\sim\mathrm{Bin}(N,(1/t)^{t})$ and $Np=B$, so
$\mathbb{P}\!\left[X\geq B\right]\geq 0.5$
by Lemma˜G.3.
The structural claims from Theorem˜1.4 remain unchanged:
Therefore,
and hence
Competitive Ratio as a function of d and B.
From (1), for all sufficiently large $d$,
$t\;\geq\;\frac{1}{2}\left(\frac{d}{2B\ln d}\right)^{\!\frac{1}{B+1}},$
and thus every online algorithm has competitive ratio at least
$\Omega\!\left(\left(\frac{d}{\ln d}\right)^{\!\frac{1}{B+1}}\right).$
@@ 8 Conclusion
We introduced a unified approach for online resource allocation with strong complementarities via static and anonymous bundle-pricing. Our mechanisms are guided by an ex-ante LP relaxation with scaled down capacities, and translate the fractional solution’s threshold structure into posted bundle prices. This yields improved guarantees as multiplicity grows: an $O(d^{1/B})$-competitive mechanism for the $d$-single-minded setting, and an $O(m^{1/(B+1)})$-competitive mechanism for general single-minded combinatorial auctions and for capacitated graph routing. We also prove information-theoretic lower bounds via an embedding of the classical prophet-inequality hard instance using qualitatively-independent partitions, showing that the dependence on $m$ (or $d$) and $B$ is essentially unavoidable. Two directions seem particularly interesting. First, closing the remaining gaps for both the general and $d$-single-minded setting. Second, sample-based implementations: our mechanism assumes that the value distributions of the buyers are explicitly given to solve the ex-ante LP and set prices. Can one obtain comparable guarantees using few samples?
@@ Appendix A Different Distribution Supports
Our results extend to the setting where each buyer $b$ has a distribution $\mathcal{D}_{b}$ over its own support $\mathcal{S}_{b}$ and by a standard scaling argument we can assume that $\mathcal{S}_{b}\subseteq\mathbb{Z}_{\geq 0}$ without loss of generality. We extend each buyer-specific support $\mathcal{S}_{b}$ to a common support $\mathcal{S}$, run our $c$-approximation algorithm in the resulting same-support instance, and post the induced randomized menu. We show that with the appropriate extension, this (randomized) menu yields a $4c/3$-approximation for the original problem with different supports.
Let $v_{\max}^{(b)}=\max_{v\in\mathcal{S}_{b}}v$ and $v_{\max}=\max_{b}v_{\max}^{(b)}$. Define $\mathcal{S}=\{0,\ldots,v_{\max}\}\supseteq\cup_{b}\mathcal{S}_{b}$. To reduce to the equal-support case, we modify each distribution $\mathcal{D}_{b}$ so that all of them share the same support $\mathcal{S}$. For each buyer $b$ with $\mathcal{S}\setminus\mathcal{S}_{b}\neq\emptyset$, define the modified distribution $\mathcal{D}_{b}^{\prime}$ by
$q^{\prime}_{b,v^{(b)}_{\max}}\;=\;q_{b,v^{(b)}_{\max}}-\varepsilon\cdot|\mathcal{S}\setminus\mathcal{S}_{b}|,\qquad q^{\prime}_{b,v}=\varepsilon\quad\forall v\in\mathcal{S}\setminus\mathcal{S}_{b},$
@@
$0<q^{\prime}_{b,v^{(b)}_{\max}}\quad\text{i.e. }\quad\varepsilon\leq C_{1}\coloneqq\min_{b}\frac{q_{b,v^{(b)}_{\max}}}{2\,|\mathcal{S}\setminus\mathcal{S}_{b}|}.$
The modified distribution $\mathcal{D}_{b}^{\prime}$ is a valid distribution, and all $\mathcal{D}_{b}^{\prime}$ share common support $\mathcal{S}$. Define $\mathcal{D}\coloneqq\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{n}$ and $\mathcal{D^{\prime}}\coloneqq\mathcal{D}^{\prime}_{1}\times\ldots\times\mathcal{D}^{\prime}_{n}$.
@@
$\min\left\{x_{b,v^{(b)}_{\max}},\varepsilon|\mathcal{S}\setminus\mathcal{S}_{b}|\right\}$
for each buyer $b$; the resulting solution is feasible for $\mathcal{D}^{\prime}$ and loses at most $\varepsilon|\mathcal{S}\setminus\mathcal{S}_{b}|\,v_{\max}^{(b)}$ in value for each buyer $b$. Thus,
$\textnormal{{FracOpt}}^{\prime}\geq\textnormal{{FracOpt}}\;-\;\varepsilon\sum_{b}|\mathcal{S}\setminus\mathcal{S}_{b}|\,v^{(b)}_{\max}$.
@@
$\varepsilon\leq C_{2}\coloneqq\min\!\left\{C_{1},\;\frac{\textnormal{{FracOpt}}}{2\sum_{b}|\mathcal{S}\setminus\mathcal{S}_{b}|\,v^{(b)}_{\max}}\right\}$,
@@
Let $I\sim\mathcal{D}$ and $I^{\prime}\sim\mathcal{D}^{\prime}$ denote value realizations under the original and modified distributions. For the total variation distance between $\mathcal{D}_{b},\mathcal{D}_{b}^{\prime}$ for each buyer $b$ we get by definition that $d_{\mathrm{TV}}(\mathcal{D}_{b},\mathcal{D}_{b}^{\prime})\;=\;\varepsilon\,|\mathcal{S}\setminus\mathcal{S}_{b}|$. Thus, since $\mathcal{D},\mathcal{D}^{\prime}$ are product distributions we get
$d_{\mathrm{TV}}(\mathcal{D},\mathcal{D}^{\prime})\;\leq\;\sum_{b}d_{\mathrm{TV}}(\mathcal{D}_{b},\mathcal{D}_{b}^{\prime})\;=\;\varepsilon\sum_{b}|\mathcal{S}\setminus\mathcal{S}_{b}|.$
@@
$0\;\leq\;\textnormal{{Alg}}_{M,I}\;\leq\;\sum_{b\in\mathcal{N}}v_{\max}^{(b)}\;\eqqcolon\;V_{\max}^{\mathrm{tot}}.$
Therefore, for any fixed $M$, using a standard property of the total variation distance (see Proposition 4.5 in [levin_peres_wilmer_markovmixing]), we get
$\Bigl|\mathbb{E}_{I}\left[\textnormal{{Alg}}_{M,I}\right]-\mathbb{E}_{I^{\prime}}\left[\textnormal{{Alg}}_{M,I^{\prime}}\right]\Bigr|\;\leq\;d_{\mathrm{TV}}(\mathcal{D},\mathcal{D}^{\prime})\cdot V_{\max}^{\mathrm{tot}}\;\leq\;\varepsilon\!\left(\sum_{b}|\mathcal{S}\setminus\mathcal{S}_{b}|\right)\!V_{\max}^{\mathrm{tot}}.$
@@
$\Bigl|\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]-\mathbb{E}_{M,I^{\prime}}\left[\textnormal{{Alg}}_{M,I^{\prime}}\right]\Bigr|\;\leq\;\varepsilon\!\left(\sum_{b}|\mathcal{S}\setminus\mathcal{S}_{b}|\right)\!V_{\max}^{\mathrm{tot}}.$
@@
$\varepsilon\;\leq\;C_{3}\coloneqq\min\left\{C_{2},\frac{\textnormal{{FracOpt}}}{4c\,\left(\sum_{b}|\mathcal{S}\setminus\mathcal{S}_{b}|\right)\!V_{\max}^{\mathrm{tot}}}\right\},$
@@
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\;\geq\;\frac{3}{4c}\,\textnormal{{FracOpt}}.$
Hence, a $c$-approximation algorithm for the equal-support case yields a $4c/3$-approximation algorithm for the general case, thus we can simply run our algorithm on the modified distributions $\mathcal{D}_{b}^{\prime}$ and post the resulting randomized menu, justifying our assumption.

### v2

@@ 10 Lower Bounds
There are $m$ items and exactly $B$ copies of each item. Let $t=t_{m,B}$ and $N:=B\,t^{t}$. By Lemma 10.3, there exists a $(B+1)$-way QI family $\{\mathcal{P}^{(1)},\dots,\mathcal{P}^{(N)}\}$ of $t$-partitions of $[m]$, where $\mathcal{P}^{(i)}=\{P^{(i)}_{1},\dots,P^{(i)}_{t}\}$. For each $(i,a)\in[N]\times[t]$, we introduce a single-minded buyer $b_{i,a}$ who is interested in $P^{(i)}_{a}$ and has value $1$ with probability $1/t$ (and 0 otherwise), independently over all $(i,a)$. We let $E_{i}$ denote the event that all $t$ buyers in group $i$ have value $1$.
The following two observations follow from the definition of the lower bound instance.
Observation 10.4.
In any feasible allocation (respecting $B$ copies per item), it is impossible to serve buyers from more than $B$ distinct groups.
Observation 10.5.
If the events $E_{i_{1}},\dots,E_{i_{B}}$ occur for $B$ distinct groups, then there exists a feasible allocation that serves all $tB$ buyers in these groups.
@@
We recall that $E_{i}$ denotes the event that all $t$ buyers in group $i$ have value $1$. Then, $\mathbb{P}\!\left[E_{i}\right]=p=(1/t)^{t}$. For $X:=\sum_{i=1}^{N}\mathbbm{1}{\left\{E_{i}\right\}}$, we have that $X\sim\mathrm{Bin}(N,p)$ and that $\mathbb{E}[X]=Np=B$. From the second case of the proof of Lemma F.4, we get that $\mathbb{P}\!\left[X\geq B\right]\geq 0.5$. ∎
Offline versus Online.
If $X\geq B$, then by Observation 10.5, the social welfare of the offline optimum is BtB\,t. Hence, the expected social welfare of the prophet is at least 0.5Bt0.5\,B\,t. By Observation 10.4, any online algorithm can collect value from at most $B$ groups. Specifically, for every group $i$, the number of buyers with positive value is $\mathrm{Bin}(t,1/t)$. Hence, once the algorithm commits to accepting buyers from some group, the expected remaining value to accept from that group is at most $1$. Thus, $\mathbb{E}[\textnormal{{Alg}}]\leq 2B$, which implies a competitive ratio of at least $t/4$. For the proof of Theorem 2.4, we use
$t=\left\lfloor\left(\frac{m}{2B\ln m}\right)^{1/(B+2)}\right\rfloor\geq\frac{1}{2}\left(\frac{m}{2B\ln m}\right)^{1/(B+2)}\,.$
Observing that $(2B)^{1/(B+2)}\leq 2$, for all $B\geq 1$, we get that $4t\geq(m/\ln m)^{1/(B+2)}$. Therefore, we obtain the following lower bound on the competitive ratio of any online algorithm, thus concluding the proof Theorem 2.4:
$\frac{1}{16}\left(\frac{m}{\ln m}\right)^{1/(B+2)}\,.$
10.2 Proof Sketch of Theorem 2.5
The proof is similar to the proof of Theorem 2.4. The only new ingredient is enforcing $|S_{b}|\leq d$. We do so by replacing Poljak’s partitions with $d$-balanced partitions (all classes have size at most $d$), obtained by sampling balanced labelings (with each label appearing exactly $d=n/t$ times). Using negative association, we get the balanced analogue $g_{\mathrm{bal}}(n,t,r,d)\ \geq\ er/t\cdot\exp\!\left(\frac{n}{rt^{r}}\right)$.
Setting $r=B+1$, $n=m=dt$ and $t=\lfloor\big(d/({2B\ln d})\big)^{1/(B+1)}\rfloor$ gives a $(B+1)$-way QI family of size $N=B\,t^{t}$ of $t$-partitions of the $m$ items with class size at most $d$. The buyers are defined exactly as before (we have one buyer per class, their value is $1$ with probability $1/t$). So, every desired bundle has size at most $d$. The binomial argument and the two structural observations remain valid and imply a lower bound of $\Omega(t)=\Omega((d/\ln d)^{1/(B+1)})$ on the competitive ratio of any online algorithm. A detailed proof of Theorem 2.5 can be found in Appendix E. ∎
@@ 11 Conclusion
We introduced a unified approach for online resource allocation with strong complementarities via static and anonymous bundle-pricing. Our mechanisms are guided by an ex-ante LP relaxation with scaled down capacities, and translate the fractional solution’s threshold structure into posted bundle prices. This yields improved guarantees as multiplicity grows: an $O(d^{1/B})$-competitive mechanism for $d$-single-minded instances, and $O(m^{1/(B+1)})$-competitiveness for general single-minded combinatorial auctions and for online network routing. We also prove information-theoretic lower bounds via an embedding of the classical prophet-inequality hard instance using qualitative-independence constructions, showing that a similar dependence on $m$ (or $d$) and $B$ is essentially unavoidable. Two directions seem particularly interesting: (i) closing the remaining gaps for both the general and the $d$-single-minded setting; and (ii) studying sample-based implementations, where the value distribution of the buyers are implicitly inferred from a small (polynomially many to begin with, ideally a constant) number of samples.
Appendix
@@ Appendix A Different Distribution Supports
Our results extend to the setting where each buyer $b$ has a distribution $\mathcal{D}_{b}$ over its own support $\mathcal{U}_{b}$ and by a standard scaling argument we can assume that $\mathcal{U}_{b}\subseteq\mathbb{Z}_{\geq 0}$ without loss of generality. We extend each buyer-specific support $\mathcal{U}_{b}$ to a common support $\mathcal{U}$, run our $c$-approximation algorithm in the resulting same-support instance, and post the induced randomized menu. We show that with the appropriate extension, this (randomized) menu yields a $4c$-approximation for the original problem with different supports.
Let $v_{\max}^{(b)}=\max_{v\in\mathcal{U}_{b}}v$ and $v_{\max}=\max_{b}v_{\max}^{(b)}$. Define $\mathcal{U}=\{0,\ldots,v_{\max}\}\supseteq\cup_{b}\mathcal{U}_{b}$. To reduce to the equal-support case, we modify each distribution $\mathcal{D}_{b}$ so that all of them share the same support $\mathcal{U}$. For each buyer $b$ with $\mathcal{U}\setminus\mathcal{U}_{b}\neq\emptyset$, define the modified distribution $\mathcal{D}_{b}^{\prime}$ by
$q^{\prime}_{b,v^{(b)}_{\max}}\;=\;q_{b,v^{(b)}_{\max}}-\varepsilon\cdot|\mathcal{U}\setminus\mathcal{U}_{b}|,\qquad q^{\prime}_{b,v}=\varepsilon\quad\forall v\in\mathcal{U}\setminus\mathcal{U}_{b},$
@@
$0<q^{\prime}_{b,v^{(b)}_{\max}}\quad\text{i.e. }\quad\varepsilon\leq C_{1}\coloneqq\min_{b}\frac{q_{b,v^{(b)}_{\max}}}{2\,|\mathcal{U}\setminus\mathcal{U}_{b}|}.$
The modified distribution $\mathcal{D}_{b}^{\prime}$ is a valid distribution, and all $\mathcal{D}_{b}^{\prime}$ share common support $\mathcal{U}$. Define $\mathcal{D}\coloneqq\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{n}$ and $\mathcal{D^{\prime}}\coloneqq\mathcal{D}^{\prime}_{1}\times\ldots\times\mathcal{D}^{\prime}_{n}$.
@@
$\min\left\{x_{b,v^{(b)}_{\max}},\varepsilon|\mathcal{U}\setminus\mathcal{U}_{b}|\right\}$
for each buyer $b$; the resulting solution is feasible for $\mathcal{D}^{\prime}$ and loses at most $\varepsilon|\mathcal{U}\setminus\mathcal{U}_{b}|\,v_{\max}^{(b)}$ in value for each buyer $b$. Thus,
$\textnormal{{FracOpt}}^{\prime}\geq\textnormal{{FracOpt}}\;-\;\varepsilon\sum_{b}|\mathcal{U}\setminus\mathcal{U}_{b}|\,v^{(b)}_{\max}$.
@@
$\varepsilon\leq C_{2}\coloneqq\min\!\left\{C_{1},\;\frac{\textnormal{{FracOpt}}}{2\sum_{b}|\mathcal{U}\setminus\mathcal{U}_{b}|\,v^{(b)}_{\max}}\right\}$,
@@
Let $I\sim\mathcal{D}$ and $I^{\prime}\sim\mathcal{D}^{\prime}$ denote value realizations under the original and modified distributions. For the total variation distance between $\mathcal{D}_{b},\mathcal{D}_{b}^{\prime}$ for each buyer $b$ we get by definition that $d_{\mathrm{TV}}(\mathcal{D}_{b},\mathcal{D}_{b}^{\prime})\;=\;\varepsilon\,|\mathcal{U}\setminus\mathcal{U}_{b}|$. Thus, since $\mathcal{D},\mathcal{D}^{\prime}$ are product distributions we get
$d_{\mathrm{TV}}(\mathcal{D},\mathcal{D}^{\prime})\;\leq\;\sum_{b}d_{\mathrm{TV}}(\mathcal{D}_{b},\mathcal{D}_{b}^{\prime})\;=\;\varepsilon\sum_{b}|\mathcal{U}\setminus\mathcal{U}_{b}|.$
@@
$0\;\leq\;\textnormal{{Alg}}_{M,I}\;\leq\;n\cdot v_{\max}\;\eqqcolon\;V_{\max}^{\mathrm{tot}}.$
Therefore, for any fixed $M$, using a standard property of the total variation distance (see Proposition 4.5 in Levin et al. [2009]), we get
$\Bigl|\mathbb{E}_{I}\left[\textnormal{{Alg}}_{M,I}\right]-\mathbb{E}_{I^{\prime}}\left[\textnormal{{Alg}}_{M,I^{\prime}}\right]\Bigr|\;\leq\;d_{\mathrm{TV}}(\mathcal{D},\mathcal{D}^{\prime})\cdot V_{\max}^{\mathrm{tot}}\;\leq\;\varepsilon\!\left(\sum_{b}|\mathcal{U}\setminus\mathcal{U}_{b}|\right)\!V_{\max}^{\mathrm{tot}}.$
@@
$\Bigl|\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]-\mathbb{E}_{M,I^{\prime}}\left[\textnormal{{Alg}}_{M,I^{\prime}}\right]\Bigr|\;\leq\;\varepsilon\!\left(\sum_{b}|\mathcal{U}\setminus\mathcal{U}_{b}|\right)\!V_{\max}^{\mathrm{tot}}.$
@@
$\varepsilon\;\leq\;C_{3}\coloneqq\min\left\{C_{2},\frac{\textnormal{{FracOpt}}}{4c\,\left(\sum_{b}|\mathcal{U}\setminus\mathcal{U}_{b}|\right)\!V_{\max}^{\mathrm{tot}}}\right\},$
@@
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\;\geq\;\frac{1}{4c}\,\textnormal{{FracOpt}}.$
Hence, a $c$-approximation algorithm for the equal-support case yields a $4c$-approximation algorithm for the general case, thus we can simply run our algorithm on the modified distributions $\mathcal{D}_{b}^{\prime}$ and post the resulting randomized menu, justifying our assumption.